### 18 Dec 2023 v1

@@ 4 Pipeline for quantified analysis of seizure video recordings
In addressing the complexities of vision-based semiology analysis for seizure video recordings, our proposed pipeline introduces a comprehensive methodology comprising distinct modules connected to form an integrated system. This system is unique and provides important supplementary and unbiased data to assess video-recorded seizures through clinical implementation. The proposed ecosystem encompasses key components such as motion capture of keypoints and regions of interest, facilitating the precise extraction of relevant spatial information. Further, motion signatures and time-evolving frequencies of stereotyped movements contribute invaluable insights into patterns that characterize seizure-related behaviors. The pipeline extends to the analysis of seizure disorders and robust seizure detection mechanisms. Multi-stream frameworks have been incorporated for refined seizure classification, ensuring an understanding of varied seizure types. Moreover, an inspection of aberrant behaviors, model prediction verification, and model interpretation modules enhance the interpretability and reliability of the entire system. Such a dynamic and modular pipeline is depicted in Fig. 4. A recent study [113] outlined the anticipated workflow of a “complete system” for video analysis of seizure event classification. However, this non-technical discussion overlooked crucial stages described above. The subsequent subsections provide detailed explanations of each module, elucidating their roles and contributions within this unified framework.
Figure 4: Overview of an automatic seizure video evaluation system. A potential implementation could be a closed-loop system designed for clinical deployment in the hospital that integrates: A. A flexible and modular markerless motion capture solution, B. A method enabling the inspection of the stepwise progression of clinical features and the order of signs, C. An analysis of the whole body to distinguish seizure disorders such as PNES, D. A flexible and modular multi-stream framework to evaluate isolated semiologies, E. An inspection of the classification results and model uncertainty (model confidence), as well as a set of model-agnostic methods to interpret deep learning methodologies, and F. An anomaly detection module to update the system with aberrant video features of semiology.
@@ 4.2 Seizure detection and temporal analysis
A comprehensive examination of the evolution of seizure-related clinical manifestations in patients with epilepsy is imperative. From an algorithmic perspective, the initial phase involves the detection or identification of a relevant time range for further analysis. The generic model performing this task must be versatile enough to accommodate various seizure types, allowing algorithms to isolate moments of interest and compare extracted features to predefined thresholds. Diverse research and commercial options were documented in Section 3.1. The analysis of seizure evolution, aimed at discerning the presence or absence of specific movement features and dynamic changes, is a major component of epilepsy patient assessment. Seizure manifestations that can be readily detected in present video analyses encompass a spectrum, ranging from repetitive rhythmic movement of trunks, limbs or hands, such as whole body rocking or manipulation of an object; to a more intense presentation characterized by excessive amounts of amplitude, speed, and acceleration [32, 28]. As well as body/limb movement patterns, seizure-related facial changed may bring valuable semiologic information, e.g., muscular contraction, eye deviation or emotional facial expression. This module, illustrated in Figure 4, focuses on the development of a computer-based tool designed to analyze the stepwise progression of clinical features and the propagation of semiology. This tool has the capability to discern dominant features, categorizing them based on speed, periodicity, the frequency of events, and differences in motions. Moreover, the module explores the correlation of motion across different body parts, given that epileptic seizures often transition from one clinical manifestation to another. The incorporation of a motion signature or semiological fingerprint proves to be valuable when scrutinizing videos, presenting semiology as a flow of signs [11]. To enhance the interpretability of the identified kinematic data, human experts can visually assess the time-evolving frequency properties of seizures [90] which may support an understanding of the mechanisms underlying such behaviors. Additional details regarding an automated fine-grained action analysis are discussed in Section 5.3.
4.3 Analysis of seizure disorders and epilepsy detection
@@
Prior to the execution of vision-based semiology classification for epilepsies, the “Controller” in the system will exclusively provide to the “Process” module seizure disorders categorized as epileptic seizures. As discussed previously in Section 3.1, seizure semiology is an essential component for patient monitoring and can differentiate between epileptic and non-epileptic seizures. This module focuses on identifying PNES or FND, which lack the characteristic changes in cortical activity associated with epilepsy, leading to misdiagnosis. To address this, the module incorporates end-to-end frameworks that analyze the entire body simultaneously, adopting proper seizure-level prediction [162] capturing semiological features across the entirety of arbitrarily long seizure, rather than averaging all snippet-level predictions from short clips. Averaging predictions is effective when the same action occurs along most of the seizure video duration, but given the variable propagation of semiology, this is not the typical case. This module utilizes temporal segment networks to split videos of any duration into $n$ non-overlapping segments and extract spatial features from input frames [162], as opposed to traditional CNNs [4]. Subsequently, the module leverages transformers to effectively learn temporal relations between the extracted spatial features of the seizure patterns, as demonstrated in previous works [88, 141]. The transformer-based pre-training approach can learn robust and generalizable features, taking advantage of large unannotated data from medical databases of past seizure records. To ensure generalization when annotated examples are scarce, the module can be flexible by incorporating emerging approaches such as meta-learning [136] and domain adaptation [55] which are becoming particularly valuable in the medical domain.
@@ 4.4 Multi-stream frameworks for seizure classification
After the presence of an epileptic seizure is confirmed by the preceding module, the focus of the “Process” module shifts to the analysis of diverse clinical manifestations, including isolated semiologies captured from one or more cameras. The models employed in this module are designed to serve as detectors for common lateralization and localization signs, including phenomena such as head version, mouth automatism, dystonic limb posturing, hand automatisms [18], and even specific signs like the thumbs-up gesture [202]. The multi-stream framework exhibits flexibility in capturing dynamic changes through appearance, keypoints, and the analysis of the 3D distribution of body parts, face and hand landmarks, employing either a fusion approach [10, 89] or a hierarchical approach [5]. Depending on the specific downstream task, such as classifying GTCS, MTLE, ETLE, etc, users have the ability to identify the most relevant motion features deserving attention for a given patient, allowing customization of the analysis for the complete body [106, 88], face [9, 89], or hands [160]. Recognizing that a single semiologic sign in isolation may not provide sufficient insight, the integration of validation accuracies for each clinical manifestation is calculated. This consolidated information supports clinical experts in making the final decision. Furthermore, the observed ictal behavior can be correlated from a semi-automated perspective with neuronal activity findings, whose concordance is necessary for localization purposes, as has been demonstrated by preliminary research in [90, 233]. Methodologies discussed in this module can be further enhanced with open-set action recognition strategies [21] that require the network to correctly classify in-distribution samples and identify out-of-distribution samples. This is discussed further in Section 5.4.
@@ 4.6 Anomaly detection - aberrant epileptic seizure
Recognizing actions in the context of epilepsy is challenging due to their subjective and dynamic nature. Different clinical manifestations, such as oral and manual automatism, can have a large amount of variation in their execution, making it difficult for recognition models to generalize effectively. While a larger labeled or unlabeled semiology database could potentially enhance accuracy, especially in situations where technology and clinical expertise are available, we aim to augment our system with a “Feedback” module. This module is designed to detect unusual or aberrant behaviors, contributing to the overall robustness and adaptability of the overall framework. In addressing this task, epileptic seizures are grouped into a best-fit model using a template of stereotypical behaviors commonly observed in various forms of epilepsy, organized into libraries. This allows system to identify whether the test semiology can fit into within a set of known categories [6]. Semi-supervised video anomaly detection techniques, such as deep autoencoders, also prove effective in this context, leveraging the full potential of normal data while addressing issues of weak labeling and data imbalance [147]. However, the selection of the most suitable deep-learning method for video anomaly detection depends on the available datasets.
@@ 4.7 Future implementation
In the future, we can imagine such a system being incorporated into the Epilepsy Monitoring Unit video set-up such that clinicians can both review the video-EEG data in the manner they currently do, and benefit from a user-friendly interface that would allow additional objective automated video analysis to complement their own visual assessment (similar to the current clinical use of automated EEG or MRI quantification software). Such a system could, for example, automatically measure the frequency of seizure-related rhythmic movements, compare the similarity index between different seizures from the same patient, compute a predictive score for diagnosis (e.g. ES versus PNES), or likely cerebral localization and/or lateralization, especially for difficult semiologies (e.g., for hyperkinetic seizures). Self-learning systems based on deep learning approaches will become more accurate as data accumulates, progressively allowing for more precise categorization of different semiologic patterns, including complex ones that are currently challenging for clinicians to interpret. As privacy-preservation methods become more advanced and methods start to be used on a larger scale with more collaboration and data sharing between clinical teams, knowledge of neural correlates of different semiologic patterns will expand. Knowledge sharing will likely be helped by similar quantified methods being applied to simultaneous electroencephalographic data (especially stereo-electroencephalography), overseen by expert clinicians and using seizure freedom after epilepsy surgery as an additional ground truth where appropriate. Once validated in specialist centres, the developed techniques may then be applicable to increasingly apply automated video analysis of seizures in ambulatory settings, maximizing access to accurate seizure diagnosis and meeting a critical current gap in epilepsy care. Beyond the epilepsy monitoring unit environment, other seizure video recording methods including smartphones might benefit from some inbuilt automated tools to complement the clinicians’ visual analysis, which could make precise and timely diagnosis of seizures more widely available.
@@ 5.1 Clinical challenges of motion capture and domain adaptation
Motion analysis is one of the main approaches that can assist medical practitioners in analyzing semiology. In [103], authors reviewed MoCap methods in relation to medical challenges including occlusions, viewpoints variations, and low-resolution 3D MoCap; and suggested paradigms that can be considered to address these challenges. Here, we provide some outline MoCap challenges and guidelines that can be leveraged to assist with semiology analysis. The repertoire of movement patterns and facial expressions observed in human seizures is extremely large, complex and heterogeneous compared to those studied in animal models, thus ideally suited to benefit from machine learning approaches but also very challenging in terms of obtaining adequate test data sets with expert clinician labelling.
@@
Advancements in computer vision MoCap methods have demonstrated high-performance gains, particularly on public datasets. For example, in domains like 3D face reconstruction and 3D pose estimation, novel techniques such as transformers or graph-based methods have been introduced. These methods show promise in addressing challenges such as occlusion and providing a more robust representation of dynamic features in faces, hands, or the body. The advancements, which have shown success in various applications, warrant exploration in the context of seizure semiology analysis.
@@ 5.2 Preserving privacy in patient monitoring for epilepsy
Visual techniques, also known as de-identification techniques, are designed to enhance visual privacy. The prevalent approach in semiology data involves transforming the image or video data into another modality, such as converting video to optical flow. These pipelines [75, 214, 73, 141] can operate locally within the hospital, safeguarding patient privacy while providing motion semiotics of seizures. However, they fall short in providing context for the analysis of specific behaviors of interest, such as small facial or hand semiology. For example, facial seizure manifestations (i.e., involuntary seizure-induced gaze deviation, specific facial contraction patterns, or emotional facial expression) may bring key information to clinical seizure analysis are often not visible in optical flow images.
Another visual-based approach is image swapping/synthesis, which is employed to conceal a patient’s identity while conveying information not easily leveraged by optical flow-based approaches. For instance,  [246] developed a face-swapping method for patients with Parkinson disease, and demonstrated that face swapping could preserve the facial landmarks while hiding the patient’s identity. This approach was also investigated for semiology  [86]. Swapping the whole body with an avatar can also be beneficial, and such techniques are based on estimating landmarks/pose, generating a high-quality mesh, and replacing the human body in a video or image with an avatar [171, 156]. Image or video synthesis can also be used at a higher level to generate synthetic data for semiology. For instance, advanced methods that take text [194] or physics-based approaches [231] along with semiology expert knowledge can be investigated to generate such data.
@@
Federated learning has gained widespread attention for its capacity to train machine learning models while preserving privacy. Positioned as a system-level privacy mechanism, federated learning ensures the protection of patients’ privacy by keeping their information localized, eliminating the need for data interchange. This decentralized approach is designed to learn from heterogeneous datasets, converging into a global model that upholds local privacy standards khan2021federated,khokhar2022review. Li et al. [121] proposed a federated learning approach to train action recognition models without central collection of users’ sensor data.
@@ 5.3 Fine-grained analysis of actions for the propagation of semiology
Epileptic seizures often transition from one clinical manifestation to another, reflecting the anatomical spread of the electrical seizure activity, which reflects the respective affected brain regions. Knowledge of clinically plausible seizure propagation pathways can therefore make a valuable contribution to the differentiation of focal seizure types, their differential diagnoses and the localization of seizure activity [149]. Existing action recognition models, as discussed in Section 3, often focus on recognizing full video segments without capturing subtle transitions or dependencies between different clinical manifestations. This limitation hinders the understanding of the temporal evolution of seizures from onset to termination. The typical propagation pathways of seizure activity determine the serial occurrence of semiological features, offering valuable supplementary information.

### 04 Mar 2024 v2

@@ 5 Pipeline for quantified analysis of seizure video recordings
In addressing the complexities of vision-based semiology analysis for seizure video recordings, our proposed pipeline introduces a comprehensive methodology comprising distinct modules connected to form an integrated system. This system is unique and provides important supplementary and unbiased data to assess video-recorded seizures through clinical implementation. The proposed ecosystem encompasses key components such as motion capture of keypoints and regions of interest, facilitating the precise extraction of relevant spatial information. Further, motion signatures and time-evolving frequencies of stereotyped movements contribute invaluable insights into patterns that characterize seizure-related behaviors. The pipeline extends to the analysis of seizure disorders and robust seizure detection mechanisms. Multi-stream frameworks have been incorporated for refined seizure classification, ensuring an understanding of varied seizure types. Moreover, an inspection of aberrant behaviors, model prediction verification, and model interpretation modules enhance the interpretability and reliability of the entire system. Such a dynamic and modular pipeline is depicted in Figure 4. The subsequent subsections provide detailed explanations of each module, elucidating their roles and contributions within this unified framework.
Figure 4: Overview of an automatic seizure video evaluation system. A potential implementation could be a closed-loop system designed for clinical deployment in the hospital that integrates: A. A flexible and modular markerless motion capture solution, B. A method enabling the inspection of the stepwise progression of clinical features and the order of signs, C. An analysis of the whole body to distinguish seizure disorders such as PNES, D. A flexible and modular multi-stream framework to evaluate isolated semiologies, E. An inspection of the classification results and model uncertainty (model confidence), as well as a set of model-agnostic methods to interpret deep learning methodologies, and F. An anomaly detection module to update the system with aberrant video features of semiology. ES: Epileptic seizures; FND: functional neurological disorder; PNES: psychogenic non-epileptic seizures.
@@ 5.2 Seizure detection and temporal analysis
A comprehensive examination of the evolution of video-recorded seizure-related clinical manifestations in patients with Epilepsy is imperative. From an algorithmic perspective, the initial phase involves the detection or identification of a relevant time range for further analysis. The generic model performing this task must be versatile enough to accommodate various seizure types, allowing algorithms to isolate moments of interest and compare extracted features to predefined thresholds. Diverse research and commercial options were documented in Section 4.1. The analysis of seizure evolution, aimed at discerning the presence or absence of specific movement features and dynamic changes, is a major component of Epilepsy patient assessment. Seizure manifestations that can be readily detected in present video analyses encompass a spectrum, ranging from repetitive rhythmic movement of trunks, limbs or hands, such as whole body rocking or manipulation of an object; to a more intense presentation characterized by excessive amounts of amplitude, speed, and acceleration [35, 29]. As well as body/limb movement patterns, seizure-related facial changes may bring valuable semiologic information, e.g. muscular contraction, eye deviation or emotional facial expression [105]. This module, illustrated in Figure 4, focuses on the development of a computer-based tool designed to analyze the stepwise progression (or simultaneous occurrence) of clinical features. This tool has the capability to discern dominant features, categorizing them based on speed, periodicity, the frequency of events, and differences in motions. Moreover, the module explores the correlation of motion across different body parts, given that epileptic seizures often transition from one clinical manifestation to another over a certain timescale that might help predict seizure propagation. The incorporation of a motion signature or semiological fingerprint proves to be valuable when scrutinizing videos, presenting semiology as a flow of signs [11]. To enhance the interpretability of the identified kinematic data, human experts can visually assess the time-evolving frequency properties of seizures [106] which may support an understanding of the mechanisms underlying such behaviors. Additional details regarding an automated fine-grained action analysis are discussed in Section 6.3.
5.3 Analysis and detection of seizure disorders
@@
Prior to the execution of vision-based semiology classification for epilepsies, the “Controller” in the system will exclusively provide to the “Process” module seizure disorders categorized as epileptic seizures. As discussed previously in Section 4.1, seizure semiology is an essential component for patient monitoring and can differentiate between epileptic and non-epileptic seizures. This module focuses on identifying PNES or FND, which lack the characteristic changes in cortical activity associated with epilepsy, sometimes contributing to misdiagnosis. To address this, the module incorporates end-to-end frameworks that analyze the entire body simultaneously, adopting proper seizure-level prediction [182] capturing semiological features across the entirety of arbitrarily long seizure, rather than averaging all snippet-level predictions from short clips. Averaging predictions is effective when the same action occurs along most of the seizure video duration, but given the variable propagation of semiology, this is not the typical case. This module utilizes temporal segment networks to split videos of any duration into $n$ non-overlapping segments and extract spatial features from input frames [182], as opposed to traditional CNNs [4]. Subsequently, the module leverages transformers to effectively learn temporal relations between the extracted spatial features of the seizure patterns, as demonstrated in previous works [104, 161]. The transformer-based pre-training approach can learn robust and generalizable features, taking advantage of large unannotated data from medical databases of past seizure records. To ensure generalization when annotated examples are scarce, the module can be flexible by incorporating emerging approaches such as meta-learning [156] and domain adaptation [61] which are becoming particularly valuable in the medical domain.
@@ 5.4 Multi-stream frameworks for seizure classification
After the presence of an epileptic seizure is confirmed by the preceding module, the focus of the “Process” module shifts to the analysis of diverse clinical manifestations, including isolated semiologies captured from one or more cameras. The models employed in this module are designed to serve as detectors for common lateralization and localization signs, including phenomena such as head version, mouth automatism, dystonic limb posturing, hand automatisms [18, 78, 209], and even specific signs like the thumbs-up gesture [226]. The multi-stream framework exhibits flexibility in capturing dynamic changes through appearance, keypoints, and the analysis of the 3D distribution of body parts, face and hand landmarks, employing either a fusion approach [10, 105] or a hierarchical approach [5]. Depending on the specific downstream task, such as classifying GTCS, MTLE, ETLE, etc, users have the ability to identify the most relevant motion features deserving attention for a given patient, allowing customization of the analysis for the complete body [122, 104], face [9, 105], or hands [180]. Recognizing that a single semiologic sign in isolation will likely not provide sufficient insight, the integration of validation accuracies for each clinical manifestation is calculated. This consolidated information supports clinical experts in making the final decision. Furthermore, the observed ictal behavior can be correlated from a semi-automated perspective with neuronal activity findings, whose concordance is necessary for localization purposes, as has been demonstrated by preliminary research in [106, 259]. Methodologies discussed in this module can be further enhanced with open-set action recognition strategies [21] that require the network to correctly classify in-distribution samples and identify out-of-distribution samples. This is discussed further in Section 6.4.
@@ 5.6 Anomaly detection - aberrant epileptic seizure
Recognizing actions (motor behaviors) in the context of epileptic seizures is challenging due to their subjective and dynamic nature. Different clinical manifestations, such as oral and manual automatism, can have a large amount of variation in their execution, and may sometimes be related to normal behavior, making it difficult for recognition models to generalize effectively. While a larger labeled or unlabeled semiology database could potentially enhance accuracy, especially in situations where technology and clinical expertise are available, we aim to augment our system with a “Feedback” module. This module is designed to detect unusual or aberrant behaviors, contributing to the overall robustness and adaptability of the overall framework. In addressing this task, epileptic seizures are grouped into a best-fit model using a template of stereotypical behaviors commonly observed in various forms of epileptic seizures, organized into libraries. This allows the system to identify whether the test semiology can fit into within a set of known categories [6]. Semi-supervised video anomaly detection techniques, such as deep autoencoders, also prove effective in this context, leveraging the full potential of normal data while addressing issues of weak labeling and data imbalance [167]. However, the selection of the most suitable deep-learning method for video anomaly detection depends on the available datasets.
@@ 5.7 Future implementation
In the future, we can imagine such a system being incorporated into the Epilepsy Monitoring Unit video set-up such that clinicians can both review the video-EEG data in the manner they currently do, and benefit from a user-friendly interface that would allow additional objective automated video analysis to complement their own visual assessment (similar to the current clinical use of automated EEG or MRI quantification software). Such a system could, for example, automatically measure the frequency of seizure-related rhythmic movements, compare the similarity index between different seizures from the same patient, compute a predictive score for diagnosis (e.g. ES versus PNES), or predict likely cerebral localization and/or lateralization with estimation of confidence intervals. This would be most useful for complex motor semiology (e.g. patterns seen in hyperkinetic seizures) [73].
Self-learning systems based on deep learning approaches will become more accurate as data accumulates, progressively allowing for more precise categorization of different semiologic patterns, including complex ones that are currently challenging for clinicians to interpret. As privacy-preservation methods become more advanced and methods start to be used on a larger scale with more collaboration and data sharing between clinical teams, knowledge of neural correlates of different semiologic patterns will expand. Knowledge sharing will likely be helped by similar quantified methods being applied to simultaneous electroencephalographic data (especially stereo-electroencephalography, SEEG), overseen by expert clinicians and using seizure freedom after epilepsy surgery as an additional ground truth where appropriate. Once validated in specialist centers, the developed techniques may then be applicable to increasingly apply automated video analysis of seizures in ambulatory settings as well, maximizing access to accurate seizure diagnosis and meeting a critical current gap in Epilepsy care. Beyond the Epilepsy monitoring unit environment, other seizure video recording methods including smartphones [273] might benefit from some inbuilt automated tools, with lightweight processing capacity, to complement the clinicians’ visual analysis, which could make a precise and timely diagnosis of seizures more widely available.
@@ 6.1 Clinical challenges of motion capture and domain adaptation
Motion analysis is one of the main approaches that can assist medical practitioners in analyzing semiology. In [119], authors reviewed MoCap methods in relation to medical challenges including occlusions, viewpoints variations, and low-resolution 3D MoCap; and suggested paradigms that can be considered to address these challenges. Here, we provide some outline MoCap challenges and guidelines that can be leveraged to assist with semiology analysis. The repertoire of movement patterns and facial expressions observed in human seizures is extremely large, complex and heterogeneous compared to those studied in animal models [132], thus ideally suited to benefit from machine learning approaches but also very challenging in terms of obtaining adequate test data sets with expert clinician labelling.
@@
Advancements in computer vision MoCap methods have demonstrated high-performance gains, particularly on public datasets. For example, in domains like 3D face reconstruction and 3D pose estimation, novel techniques such as transformers or graph-based methods have been introduced. These methods show promise in addressing challenges such as occlusion and providing a more robust representation of dynamic features in faces, hands, or the body. The advancements, which have shown success in various applications, warrant exploration in the context of seizure semiology analysis. As experience and sensitivity/specificity of methods improved, more subtle discrimination will be possible, e.g. distinguishing between ASD-related hand stereotypies and epileptic movements in a patient with both conditions. As such, future work may concern the prediction of depression scores from facial expression analysis etc, in patients with epilepsy, i.e. detecting presence of comorbidities in the resting state (between-seizure periods), which could be considered as “inter-ictal semiology”.
@@ 6.2 Preserving privacy in patient monitoring for Epilepsy
Visual techniques, also known as de-identification techniques, are designed to enhance visual privacy. The prevalent approach in semiology data involves transforming the image or video data into another modality, such as converting video to optical flow. These pipelines [91, 239, 84, 161] can operate locally within the hospital, safeguarding patient privacy while providing motion semiotics of seizures. However, they fall short in providing context for the analysis of specific behaviors of interest, such as small facial or hand semiology. For example, facial seizure manifestations (i.e. involuntary seizure-induced gaze deviation, specific facial contraction patterns, or emotional facial expression) may bring key information to clinical seizure analysis are often not visible in optical flow images.
Another visual-based approach is image swapping/synthesis, which is employed to conceal a patient’s identity while conveying information not easily leveraged by optical flow-based approaches. For instance,  [272] developed a face-swapping method for patients with Parkinson disease, and demonstrated that face swapping could preserve the facial landmarks while hiding the patient’s identity. The efficacy of different face transformation approaches was then investigated in patients with facial semiology captured during epileptic seizures [102]. Swapping the whole body with an avatar can also be beneficial for privacy protection, and such techniques are based on estimating landmarks/pose, generating a high-quality mesh, and replacing the human body in a video or image with an avatar [192, 176]. Image or video synthesis can also be used at a higher level to generate synthetic data for semiology. For instance, advanced methods that take text [217] or physics-based approaches [257] along with semiology expert knowledge can be investigated to generate such data.
@@
Federated learning has gained widespread attention for its capacity to train machine learning models while preserving privacy. Positioned as a system-level privacy mechanism, federated learning ensures the protection of patients’ privacy by keeping their information localized, eliminating the need for data interchange. This decentralized approach is designed to learn from heterogeneous datasets, converging into a global model that upholds local privacy standards [128, 129]. Li et al. [141] proposed a federated learning approach to train action recognition models without central collection of users’ sensor data.
@@ 6.3 Fine-grained analysis of actions for the propagation of semiology
Epileptic seizures are characterized by multiple clinical features which transition over a short timescale (seconds-minutes), reflecting the anatomical spread of the electrical seizure activity, in both its spatial (anatomic) and temporal (e.g. discharge frequency) aspects [159]. Knowledge of clinically plausible seizure types and propagation pathways can therefore make a valuable contribution to the differentiation of focal seizure types, their differential diagnoses and the cerebral organization of seizure activity [169]. Existing action recognition models, as discussed in Section 4, often focus on recognizing full video segments without capturing subtle transitions or dependencies between different clinical manifestations. This limitation hinders the understanding of the temporal evolution of seizures from onset to termination.